Calculate the wavelength of Radio 4 which broadcasts on a frequency of 198, k, H, z,


198kHz.


1
Expert's answer
2021-11-29T11:41:11-0500
"\lambda=\dfrac{c}{f}=\dfrac{3\times10^8\ \dfrac{m}{s}}{198\times10^3\ Hz}=1515\ m."